OpenStudy (anonymous):

I have a logarithm equation that I need to solve, and I don't understand how to solve it. Here is the problem: log (y x^6) = blank

OpenStudy (anonymous):

that is a mystery perhaps you are supposed to write \[\log(y)+6\log(x)\] but that is not solving anything

OpenStudy (anonymous):

it is just rewriting in another form using properties of the log

OpenStudy (anonymous):

Is this your question? \[\log_{y} x^6\]

OpenStudy (anonymous):

si

OpenStudy (anonymous):

how about this problem? x125^1/3?

OpenStudy (anonymous):

\[y^6=x\] Then that's your answer.

OpenStudy (anonymous):

@needmathhelpplease could you use the equation tab to write your question?